Is y-5=2(x-1) an equation of a line through (4,11)? Need help asap! Thank you!
Answer by rothauserc(4718)

y-5=2(x-1)
put this equation in standard line format, y = mx +b
y = 2x +3
now, is point (4, 11) on the line, substitute for x and y
11 = 2*4 +3
11 = 8 +3
11 = 11
yes, point (4, 11) is on the line